OpenStudy (anonymous):

Write the equation of a line parallel to y=9x+3 that passes through the point (2,6)

OpenStudy (anonymous):

parallel means that there slopes are the same so y= mx+ b m = slope

OpenStudy (anonymous):

using point slope formula we can find the an equation using the slope that passes through the point y-y1= m(x - x1)

OpenStudy (anonymous):

ohhhh! I remember how to do it know thanks a lot!!

OpenStudy (anonymous):

@alyssa_xo do you know java?

OpenStudy (alyssa_xo):

I do, I do.

OpenStudy (anonymous):

i don't need help lol, i just saw you post something on a java question. I don't see too many girls answering comp science questions. Are you in college?

OpenStudy (alyssa_xo):

I'm a guy pretending to be a girl, but yeah I'm studying CS in college. Big mistake given my school's CS dept is kinda dead

OpenStudy (anonymous):

oh lol nice. i was about to say most attractive cs girl i have ever seen. Good thing i didn't say that! :)

OpenStudy (anonymous):

haha lol @alyssa_xo

OpenStudy (alyssa_xo):

LOL there's this one girl that was in my intro to c++ class, haley, I believe She was better than this girl. She also started a women's support group for CS majors or something. Several of my classes have been 100% male (analysis of algorithms for one) and that's depressing especially if it's telling of the job market's gender equality, but I think it's getting better.

OpenStudy (anonymous):

what school do you go to?

OpenStudy (anonymous):

hey @Austin6i6 will you help me answer another question? its pretty similar to this one

OpenStudy (anonymous):

sure

OpenStudy (alyssa_xo):

Brooklyn College, It's part of the CUNY (City uni of NY) system. What are you in? You seem to have done a lot more math than I have. My major ended at calc 2/discrete math/linear algebra A friend in another school in the CUNY system had to do calc 3

OpenStudy (anonymous):

okay thanks its write the equation of a line perpendicular to y=3/2x-1 with a y-intercept of -14

OpenStudy (anonymous):

I attend Michigan state university, I am taking my final math class calc 3 class right now. I might have to take discrete though. I had to take calculus based physics sequences which i credit most of my math skills too. did you have to take that pellet for comp sci?

OpenStudy (anonymous):

@hl147 sorry ok now im ready

OpenStudy (anonymous):

thats fine! :)

OpenStudy (anonymous):

so if its perpendicular the slope will be the negative reciprocal.

OpenStudy (anonymous):



OpenStudy (alyssa_xo):

Thankfully, no. There a general requirement on physics 1, (or you can take geology) for a 4 year degree though. my school is a joke in terms of "seriousness", imo. sorry for derailing your question hl, austin feel free to pm me if you want to continue talking

OpenStudy (anonymous):

so on the first question i asked im not sure what to do next